OpenStudy (anonymous):

can somebody please explain the purpose of Newton’s Method. State the Newton’s Method Formula and explain how the method approximates zeros

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Newton's_method The gif on that page explains it better than I can.

OpenStudy (anonymous):

thanks can you please define the zero of a function?

OpenStudy (anonymous):

Haha oooh. Well, the zeros of a function are the inputs, or Xs, that give an output, or Y of 0. So consider the function: f(x) = 2x -6 if I input 3, the output is f(3) = 2(3) - 6 = 6-6 = 0 so 3 is the zero of the function. What this looks like on the graphs is: